OpenStudy (anonymous):

What is 5 1/5 as an improper fraction?

OpenStudy (anonymous):

\[\frac{ 26 }{ 5 }\]

OpenStudy (jagatuba):

It is very easy to convert a mixed number to an improper fraction. First you know that the denominator of your improper fraction is going to be the same as the denominator in the fraction of your mixed number so just bring that over like this:\[5\frac{ 1 }{ 5 }=\frac{ ? }{ 5 }\]Then to find the numerator of the improper fraction all you do is multiply the denominator of the fraction with the whole number and then ad the numerator. In this case the numerator of the improper fraction is 5 x 5 + 1 = 26. Or more visually:\[5\frac{ +1 }{ \times5 }=\frac{ 26 }{ 5 }\]